# On local Galois deformation rings: generalised reductive groups

### v1

@@ 1. Introduction
Let $p$ denote any prime number, let $F$ be a finite extension of $\mathbb{Q}_{p}$, and let $\Gamma_{F}$ denote its absolute Galois group. Let $L$ be another finite extension of $\mathbb{Q}_{p}$ with ring of integers $\mathcal{O}$, uniformizer $\varpi$ and residue field $k=\mathcal{O}/\varpi$.
@@
We expect that our results will play an important role in the categorical $p$-adic local Langlands correspondence. Conjecturally [27, Conjecture 6.1.14] on the Galois side one should consider a derived category of coherent sheaves (or some $\infty$-category version of it) on the Emerton–Gee stack. The local rings considered in 1.1 should arise as versal rings at finite type points of the Emerton–Gee stack, when $G$ is an $L$-group or a $C$-group of a connected reductive group defined over $F$ (see [26, Proposition 3.6.3], when $G=\operatorname{GL}_{d}$). In other words, they are expected to describe the local properties of the Emerton–Gee stack. An instance of this is a paper of Yu Min [42], where he uses the analogue of 1.1 proved in [8] for $\operatorname{GL}_{d}$ to show that if $G=\operatorname{GL}_{d}$ then the derived version of the Emerton–Gee stack is equivalent to the classical Emerton–Gee stack, obtaining an analogue of the result explained above, that derived deformation rings are homotopy discrete, over the whole of the stack. The Emerton–Gee stack has been recently defined and studied for $\mathrm{GSp}_{4}$ in [38] and for tame groups in [39]. We infinitesimally guarantee that the results of [42] will extend to this context.
In the paper we introduce and study a scheme $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$; its definition is reviewed in Section 1.4.1 below. Its $\overline{k}$-points correspond to continuous representations $\rho:\Gamma_{F}\rightarrow G(\overline{k})$ with $G$-semisimplification equal to $\overline{\rho}^{\mathrm{ss}}$. It carries an action of $G^{0}$ and the quotient stack $[X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}/G^{0}]$ is an analogue of the stacks studied by Wang-Erickson in [57] for $G=\operatorname{GL}_{d}$. We expect that the disjoint union of the stacks $[X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}/G^{0}]$ (or rather their formal versions) taken over all possible $\overline{\rho}^{\mathrm{ss}}$ will be the largest substack of the Emerton–Gee stack on which the universal $(\varphi,\Gamma)$-module can be realised as a representation of $\Gamma_{F}$, i.e. the situation described in [26, Remark 6.7.4] for $\operatorname{GL}_{d}$ will continue to hold in a more general setting.
@@ 1.1. The conjecture of Böckle–Juschka
If $\pi_{1}(G^{\prime})$ is étale then the rings $R^{\square,\chi}_{\overline{\rho}}$, $R^{\square,\chi}_{\overline{\rho}}/\varpi$ are complete intersection, regular in codimension $[F:\mathbb{Q}_{p}]$, normal domains.
@@
If $\pi_{1}(G^{\prime})$ is étale and $[F:\mathbb{Q}_{p}]\geq 3$ then $R^{\square,\chi}_{\overline{\rho}}$ and $R^{\square,\chi}_{\overline{\rho}}/\varpi$ are factorial.
If a further hypothesis on $\overline{\rho}$ and $G$ is satisfied (for example, if $G/Z(G^{0})$ does not have $\operatorname{PGL}_{2}$ as a factor) then we show in Corollary 15.20 that $R^{\square,\chi}_{\overline{\rho}}$ and $R^{\square,\chi}_{\overline{\rho}}/\varpi$ are regular in codimension $2[F:\mathbb{Q}_{p}]-1$ and hence 1.4 holds if $[F:\mathbb{Q}_{p}]=2$.
@@ 1.3. Deformation space of Lafforgue’s pseudocharacters
If $G=\operatorname{GL}_{d}$ then Emerson–Morel have shown in [25] that there is a natural bijection between $G$-pseudocharacters and Chenevier’s $d$-dimensional determinant laws defined in [16]. Thus in this case the ring $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ coincides with the deformation ring studied by Böckle–Juschka in [9]. The results of that paper are an important input in [8]. In our paper we extend the main result [9, Theorem 5.5.1] to $G$-pseudocharacters, also obtaining a new proof of it, when $G=\operatorname{GL}_{d}$.
@@ 1.4. Complete intersection
    the GIT-quotient $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}\sslash G^{0}$ is the spectrum of a complete local noetherian $\mathcal{O}$-algebra $R^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$ with residue field $k$.
@@ 1.4.1. Definition
If $G=\operatorname{GL}_{d}$ then one may linearise representations and work instead with Cayley–Hamilton homomorphisms of $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-algebras $E^{u}\rightarrow M_{d}(A)$, where $E^{u}$ is a universal Cayley–Hamilton quotient of the completed group algebra $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\llbracket\Gamma_{F}\rrbracket$. We refer the reader to Section 5.1 for more details. Wang-Erickson has shown in [57] that the algebra $E^{u}$ is a finitely generated $R^{\mathrm{ps}}$-module. This result is the reason why $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-condensed representations do not explicitly appear in [8]. In Lemma 5.2 we show that $X^{\mathrm{gen}}_{\operatorname{GL}_{d},\overline{\rho}^{\mathrm{ss}}}$ coincides with the scheme $X^{\mathrm{gen}}$ defined in [8]. In particular, it is represented by a finite type $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-algebra $A^{\mathrm{gen}}_{\operatorname{GL}_{d}}$.
@@
Let $S$ be a locally noetherian scheme and let $f:H\rightarrow G$ be a finite morphism of generalised reductive smooth affine $S$-group schemes. Then the induced map on the GIT-quotients
@@
The homomorphism $R^{\mathrm{ps}}_{G}\rightarrow A^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$ is $G^{0}$-equivariant for the trivial action on the source. Hence, its image is contained in $G^{0}$-invariants $R^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$.
@@
The morphism $R^{\mathrm{ps}}_{G}\rightarrow R^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$ is a finite universal homeomorphism.
We expect that the morphism is adequate in the sense of Alper [3], i.e. that moreover $R^{\mathrm{ps}}_{G}[1/p]\cong R^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}[1/p]$, and hope to return to this question in future work. If this were true then one could show that $\operatorname{Spec}R^{\mathrm{ps}}_{G}[1/p]$ is reduced and hence normal, as explained in Section 1.3.
@@ 1.4.2. Bounding the dimension
One is left to understand the locus, where $H^{2}(\Gamma_{F},\operatorname{ad}^{0}\rho_{x})\neq 0$. By local Tate duality this is equivalent to $H^{0}(\Gamma_{F},(\operatorname{ad}^{0}\rho_{x})^{*}(1))\neq 0$. Let us suppose that $\rho_{x}$ is absolutely $G$-irreducible. If $G=\operatorname{GL}_{d}$ then Böckle–Juschka show in [9] that non-vanishing of $H^{0}(\Gamma_{F},(\operatorname{ad}^{0}\rho_{x})^{*}(1))$ implies that $\rho_{x}$ is induced from $\Gamma_{F^{\prime}}$, where $F^{\prime}$ is a proper abelian extension of $F$. We know the dimension of $\overline{X}^{\mathrm{ps},\Gamma_{F^{\prime}}}_{\operatorname{GL}_{d^{\prime}}}$, where $d^{\prime}=\frac{d}{[F^{\prime}:F]}$, inductively and using this it is shown in [8] that the closure of this locus has positive codimension. Such arguments might work for classical groups (JQ has studied the case $G=\operatorname{Sp}_{2n}$ under the assumption $p>2$ in [44, Section 7]), but it seems impossible to get a characterisation like this for an arbitrary generalised reductive group.
@@
The main input in our proof of the above result is Matsushima’s theorem.
@@
Let us go back to the situation at hand, where $H$ is the Zariski closure of $\rho_{x}(\Gamma_{F})$ in $G(\kappa)$. If $x$ lies in the special fibre, then the twist by the cyclotomic character vanishes after restriction to $\Gamma_{F(\zeta_{p})}$. Using this we show that non-vanishing of $H^{0}(\Gamma_{F},(\operatorname{ad}\rho_{x})^{*}(1))$ implies that the neutral component $H^{0}$ of $H$ has non-zero invariants in $(\operatorname{Lie}G^{\prime})^{*}_{\kappa}$. If $\pi_{1}(G^{\prime})$ is étale or $\mathrm{char}(\kappa)=0$ then $(\operatorname{Lie}G^{\prime})^{*}_{\kappa}=(\operatorname{Lie}G^{\prime}_{\operatorname{sc}})^{*}_{\kappa}$ and Proposition 1.9 and Lemma 1.10 imply that $\dim G_{\kappa}-\dim H^{0}\geq 2$ and hence $\dim G_{\kappa}-\dim H\geq 2$. The fact that the difference can be bounded below by $2$ (and not by $1$) plays an important role in verifying Serre’s criterion for normality in the proof of 1.3. As you can see there is an induction argument shaping up. To explain it we need two further ingredients.
@@
We refer to $X_{W,j}$ as $W$-special locus of level $j$ and just $W$-special locus when $j=0$. The complement of the $W$-special locus is called $W$-non-special. The proof amounts to reformulating these conditions in terms of some finitely generated $A$-module and a little of commutative algebra. We apply this notion to $W=(\operatorname{Lie}G^{\prime}_{\operatorname{sc}})^{*}$ and if the image of $\rho$ is contained in $P(A)$, where $P$ is a parabolic subgroup scheme of $G$ with unipotent radical $U$, then we apply it to $W=(\operatorname{Lie}U)^{*}$. When working with the special fibre we consider representations $W$ over $k$ instead of $\mathcal{O}$.
@@ 1.5. Irreducible components
where $R^{\square,\chi}_{G/G^{\prime},\varphi\circ\rho_{x}}$ is a regular ring. Further, if $x$ is in the $(\operatorname{Lie}G^{\prime})^{\ast}$-non-special locus then $s$ in the above presentation is 00 and we can conclude that $R^{\square,\chi}_{G,\rho_{x}}$ is a regular ring. We deduce that the $(\operatorname{Lie}G^{\prime})^{\ast}$-non-special locus in $X^{\mathrm{gen},\chi}_{G,\overline{\rho}^{\mathrm{ss}}}$ is contained in the regular locus. In general, the $(\operatorname{Lie}G^{\prime})^{\ast}$-non-special locus might be empty, for example this happens if $G=\operatorname{PGL}_{p}$. On the other hand the complement of $(\operatorname{Lie}G^{\prime}_{\operatorname{sc}})^{\ast}$-non-special locus has codimension of at least $1+[F:\mathbb{Q}_{p}]$ in $X^{\mathrm{gen},\chi}_{G,\overline{\rho}^{\mathrm{ss}}}$. If $\pi_{1}(G^{\prime})$ is étale then these loci coincide and this implies that $X^{\mathrm{gen},\chi}_{G,\overline{\rho}^{\mathrm{ss}}}$ is regular in codimension $[F:\mathbb{Q}_{p}]$. Since $X^{\mathrm{gen},\chi}_{G,\overline{\rho}^{\mathrm{ss}}}$ is excellent the same applies to the completions of its local rings, and we may deduce that the deformation rings $R^{\square,\chi}_{G,\rho_{x}}$ are regular in codimension $[F:\mathbb{Q}_{p}]$. To show that they are complete intersection it is enough to bound their dimension from above and then use (6). Since $k[\mu]$ is a local artinian algebra the underlying reduced subschemes of the special fibres of $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$ and $X^{\mathrm{gen},\chi}_{G,\overline{\rho}^{\mathrm{ss}}}$ coincide and hence they have the same dimension. Serre’s criterion for normality implies that $R^{\square,\chi}_{G,\rho_{x}}$ is normal. Since the ring is local, it is a domain.

### v2

@@ 1. Introduction
Let $p$ denote any prime number, let $F$ be a finite extension of $\mathbb{Q}_{p}$, and let $\Gamma_{F}$ denote its absolute Galois group. Let $L$ be another finite extension of $\mathbb{Q}_{p}$ with ring of integers $\mathcal{O}$, uniformiser $\varpi$ and residue field $k=\mathcal{O}/\varpi$.
@@
We expect that our results will play an important role in the categorical $p$-adic local Langlands correspondence. Conjecturally [26, Conjecture 6.1.14] on the Galois side one should consider a derived category of coherent sheaves (or some $\infty$-category version of it) on the Emerton–Gee stack. The local rings considered in 1.1 should arise as versal rings at finite type points of the Emerton–Gee stack, when $G$ is an $L$-group or a $C$-group of a connected reductive group defined over $F$ (see [27, Proposition 3.6.3], when $G=\operatorname{GL}_{d}$). In other words, they are expected to describe the local properties of the Emerton–Gee stack. An instance of this is a paper of Yu Min [45], where he uses the analogue of 1.1 proved in [6] for $\operatorname{GL}_{d}$ to show that if $G=\operatorname{GL}_{d}$ then the derived version of the Emerton–Gee stack is equivalent to the classical Emerton–Gee stack, obtaining an analogue of the result explained above, that derived deformation rings are homotopy discrete, over the whole of the stack. The Emerton–Gee stack has been recently defined and studied for $\mathrm{GSp}_{4}$ in [39] and for tame groups in [40]. We infinitesimally guarantee that the results of [45] will extend to this context.
In the paper we introduce and study a scheme $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$; its definition is reviewed in Section 1.4.1 below. Its $\overline{k}$-points correspond to continuous representations $\rho:\Gamma_{F}\rightarrow G(\overline{k})$ with $G$-semisimplification equal to $\overline{\rho}^{\mathrm{ss}}$. It carries an action of $G^{0}$ and the quotient stack $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}/G^{0}]$ is an analogue of the stacks studied by Wang-Erickson in [62] for $G=\operatorname{GL}_{d}$. We expect that the disjoint union of the stacks $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}/G^{0}]$ (or rather their formal versions) taken over all possible $\overline{\rho}^{\mathrm{ss}}$ will be the largest substack of the Emerton–Gee stack on which the universal $(\varphi,\Gamma)$-module can be realised as a representation of $\Gamma_{F}$, i.e. the situation described in [27, Remark 6.7.4] for $\operatorname{GL}_{d}$ will continue to hold in a more general setting.
@@ 1.1. The conjecture of Böckle–Juschka
If $\pi_{1}(G^{\prime})$ is étale then the rings $R^{\square,\chi}_{\overline{\rho}}$, $R^{\square,\chi}_{\overline{\rho}}/\varpi$ are complete intersection, regular in codimension $F:\mathbb{Q}_{p}]$, normal domains.
@@
If $\pi_{1}(G^{\prime})$ is étale and $F:\mathbb{Q}_{p}]\geq 3$ then $R^{\square,\chi}_{\overline{\rho}}$ and $R^{\square,\chi}_{\overline{\rho}}/\varpi$ are factorial.
If a further hypothesis on $\overline{\rho}$ and $G$ is satisfied (for example, if $G/Z(G^{0})$ does not have $\operatorname{PGL}_{2}$ as a factor) then we show in Corollary 15.20 that $R^{\square,\chi}_{\overline{\rho}}$ and $R^{\square,\chi}_{\overline{\rho}}/\varpi$ are regular in codimension $2[F:\mathbb{Q}_{p}]-1$ and hence 1.4 holds if $F:\mathbb{Q}_{p}]=2$. If $F=\mathbb{Q}_{p}$ then [6, Remark 4.24, Corollary 4.25] provide examples, where 1.4 fails.
@@ 1.3. Deformation space of Lafforgue’s pseudocharacters
If $G=\operatorname{GL}_{d}$ then Emerson–Morel have shown in [25] that there is a natural bijection between $G$-pseudocharacters and Chenevier’s $d$-dimensional determinant laws defined in [15]. Thus in this case the ring $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ coincides with the deformation ring studied by Böckle–Juschka in [7]. The results of that paper are an important input in [6]. In our paper we extend the main result [7, Theorem 5.5.1] to $G$-pseudocharacters, also obtaining a new proof of it, when $G=\operatorname{GL}_{d}$.
@@ 1.4. Complete intersection
    the GIT quotient $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}\sslash G^{0}$ is the spectrum of a complete local noetherian $\mathcal{O}$-algebra $R^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$ with residue field $k$.
@@ 1.4.1. Definition
If $G=\operatorname{GL}_{d}$ then one may linearise representations and work instead with Cayley–Hamilton homomorphisms of $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-algebras $E^{u}\rightarrow M_{d}(A)$, where $E^{u}$ is a universal Cayley–Hamilton quotient of the completed group algebra $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\llbracket\Gamma_{F}\rrbracket$. We refer the reader to Section 5.1 for more details. Wang-Erickson has shown in [62] that the algebra $E^{u}$ is a finitely generated $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-module. This result is the reason why $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-condensed representations do not explicitly appear in [6]. In Lemma 5.2 we show that $X^{\mathrm{gen}}_{\operatorname{GL}_{d},\overline{\rho}^{\mathrm{ss}}}$ coincides with the scheme $X^{\mathrm{gen}}$ defined in [6]. In particular, it is represented by a finite type $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-algebra $A^{\mathrm{gen}}_{\operatorname{GL}_{d}}$.
@@
Let $S$ be a locally noetherian scheme and let $f:H\rightarrow G$ be a finite morphism of generalised reductive smooth affine $S$-group schemes. Then the induced map on the GIT quotients
@@
The homomorphism $R^{\mathrm{ps}}_{G}\rightarrow A^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$ is $G^{0}$-equivariant for the trivial action on the source. Hence, its image is contained in $G^{0}$-invariants $R^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$, which represents the GIT quotient $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}\sslash G^{0}$.
@@
The morphism $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}\sslash G^{0}\rightarrow\operatorname{Spec}R^{\mathrm{ps}}_{G}$ is a finite universal homeomorphism.
We expect that the morphism is adequate in the sense of Alper [1], i.e. that moreover $R^{\mathrm{ps}}_{G}[1/p]\cong R^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}[1/p]$, and hope to return to this question in future work [^1] We have proved this assertion in [47].. If this were true then one could show that $\operatorname{Spec}R^{\mathrm{ps}}_{G}[1/p]$ is reduced and hence normal, as explained in Section 1.3.
@@ 1.4.2. Bounding the dimension
One is left to understand the locus, where $H^{2}(\Gamma_{F},\operatorname{ad}^{0}\rho_{x})\neq 0$. By local Tate duality this is equivalent to $H^{0}(\Gamma_{F},(\operatorname{ad}^{0}\rho_{x})^{*}(1))\neq 0$. Let us suppose that $\rho_{x}$ is absolutely $G$-irreducible. If $G=\operatorname{GL}_{d}$ then Böckle–Juschka show in [7] that non-vanishing of $H^{0}(\Gamma_{F},(\operatorname{ad}^{0}\rho_{x})^{*}(1))$ implies that $\rho_{x}$ is induced from $\Gamma_{F^{\prime}}$, where $F^{\prime}$ is a proper abelian extension of $F$. We know the dimension of $\overline{X}^{\mathrm{ps},\Gamma_{F^{\prime}}}_{\operatorname{GL}_{d^{\prime}}}$, where $d^{\prime}=\frac{d}{[F^{\prime}:F]}$, inductively and using this it is shown in [6] that the closure of this locus has positive codimension. Such arguments might work for classical groups (JQ has studied the case $G=\operatorname{Sp}_{2n}$ under the assumption $p>2$ in [50, Section 7]), but it seems impossible to get a characterisation like this for an arbitrary generalised reductive group.
@@
Our original proof of the above result used Matsushima’s theorem. The current simpler proof was suggested by both Sean Cotner and the anonymous referee.
@@
Let us go back to the situation at hand, where $H$ is the Zariski closure of $\rho_{x}(\Gamma_{F})$ in $G(\kappa)$. If $x$ lies in the special fibre, then the twisting operation by the cyclotomic character becomes trivial after the restriction to $\Gamma_{F(\zeta_{p})}$. Using this we show that non-vanishing of $H^{0}(\Gamma_{F},(\operatorname{ad}\rho_{x})^{*}(1))$ implies that the neutral component $H^{0}$ of $H$ has non-zero invariants in $(\operatorname{Lie}G^{\prime})^{*}_{\kappa}$. If $\pi_{1}(G^{\prime})$ is étale or $\mathrm{char}(\kappa)=0$ then $(\operatorname{Lie}G^{\prime})^{*}_{\kappa}=(\operatorname{Lie}G^{\prime}_{\operatorname{sc}})^{*}_{\kappa}$ and Proposition 1.9 and Lemma 1.10 imply that $\dim G_{\kappa}-\dim H^{0}\geq 2$ and hence $\dim G_{\kappa}-\dim H\geq 2$. The fact that the difference can be bounded below by $2$ (and not by $1$) plays an important role in verifying Serre’s criterion for normality in the proof of 1.3. As you can see there is an induction argument shaping up. To explain it we need two further ingredients.
@@
We refer to $X_{W,j}$ as the $W$-special locus of level $j$ and just the $W$-special locus when $j=0$. The complement of the $W$-special locus is called $W$-non-special. The proof amounts to reformulating these conditions in terms of some finitely generated $A$-module and a little of commutative algebra. We apply this notion to $W=(\operatorname{Lie}G^{\prime}_{\operatorname{sc}})^{*}$ and if the image of $\rho$ is contained in $P(A)$, where $P$ is a parabolic subgroup scheme of $G$ with unipotent radical $U$, then we apply it to $W=(\operatorname{Lie}U)^{*}$. When working with the special fibre we consider representations $W$ over $k$ instead of $\mathcal{O}$.
@@ 1.5. Irreducible components
where $R^{\square,\chi}_{G/G^{\prime},\varphi\circ\rho_{x}}$ is a regular ring. Further, if $x$ is in the $(\operatorname{Lie}G^{\prime})^{\ast}$-non-special locus then $s$ in the above presentation is 0 and we can conclude that $R^{\square,\chi}_{G,\rho_{x}}$ is a regular ring. We deduce that the $(\operatorname{Lie}G^{\prime})^{\ast}$-non-special locus in $X^{\mathrm{gen},\chi}_{G,\overline{\rho}^{\mathrm{ss}}}$ is contained in the regular locus. In general, the $(\operatorname{Lie}G^{\prime})^{\ast}$-non-special locus might be empty, for example this happens if $G=\operatorname{PGL}_{p}$. On the other hand the complement of $(\operatorname{Lie}G^{\prime}_{\operatorname{sc}})^{\ast}$-non-special locus has codimension of at least $1+[F:\mathbb{Q}_{p}]$ in $X^{\mathrm{gen},\chi}_{G,\overline{\rho}^{\mathrm{ss}}}$. If $\pi_{1}(G^{\prime})$ is étale then these loci coincide and this implies that $X^{\mathrm{gen},\chi}_{G,\overline{\rho}^{\mathrm{ss}}}$ is regular in codimension $F:\mathbb{Q}_{p}]$. Since $X^{\mathrm{gen},\chi}_{G,\overline{\rho}^{\mathrm{ss}}}$ is excellent the same applies to the completions of its local rings, and we may deduce that the deformation rings $R^{\square,\chi}_{G,\rho_{x}}$ are regular in codimension $F:\mathbb{Q}_{p}]$. To show that they are complete intersection it is enough to bound their dimension from above and then use (6). Since $k[\mu]$ is a local artinian algebra the underlying reduced subschemes of the special fibres of $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$ and $X^{\mathrm{gen},\chi}_{G,\overline{\rho}^{\mathrm{ss}}}$ coincide and hence they have the same dimension. Serre’s criterion for normality implies that $R^{\square,\chi}_{G,\rho_{x}}$ is normal. Since the ring is local, it is a domain.